# Few-Shot Retinomorphic Vision in a Nonlinear Photonic Network Laser

### v4

@@ Abstract
With the growing prevalence of AI, demand increases for hardware that mimics the brain’s ability to extract structure from limited data. In the retina, ganglion cells detect features from sparse inputs via lateral inhibition, where neurons antagonistically suppress activity of neighbouring cells. Biological neurons exhibit diverse heterogenous nonlinear responses, linked to robust learning and strong performance in low-data regimes.
Here, we introduce a bio-inspired ‘retinomorphic’ photonic system where spatially-competing lasing modes in a network laser act as heterogeneous, inhibitively-coupled neurons - enabling few-shot classification and segmentation. This compact (150 m) silicon-compatible scheme addresses key challenges in photonic computing: physical nonlinearity and spatial footprint. We report 98.05% and 87.85% accuracy on MNIST and Fashion-MNIST, and 90.12% on BreaKHis cancer diagnosis - outperforming software CNNs including EfficientNet in few-shot and class-imbalanced regimes. We demonstrate combined segmentation and classification on the HAM10k skin lesion dataset, achieving DICE and Jaccard scores of 84.49% and 74.80%. These results establish a new class of nonlinear photonic hardware for versatile, data-efficient neuromorphic computing.
@@ Introduction
Physics-based neuromorphic computing systems[1, 12, 3, 13, 14, 15, 16] have been explored as candidates for edge-AI hardware[14], but to date demonstrations integrating the above qualities and providing software-equivalent performance on harder problems such as biomedical tasks have remained elusive.
@@
Biological systems offer a compelling blueprint. In the retina, ganglion cells perform spatially-sensitive nonlinear processing via lateral inhibition[17, 18, 19, 20], where neurons compete to fire while inhibitively suppressing activity of neighbouring cells. This mechanism enables efficient encoding of spatial structure, even under sparse or noisy input, and the ability to detect many features in parallel from multiple ganglion cells[17, 21]. The combination of excitatory dynamics - firing in response to input, and inhibitory ones - suppressing activity of nearby coupled neurons, dynamics is crucial to the strong performance of biological learning systems, but to date physics-based neuromorphic schemes typically focus solely on engineering excitatory dynamics. Addressing this is an inviting route towards taking neuromorphic performance closer to the biological systems which inspire it. In addition, the diverse heterogenous nonlinear responses across ensembles of biological neurons have been linked to robust learning[22], and creating effective high-dimensional latent spaces which can give rise to strong few-shot learning[11].
Among photonic systems, complex random lasers hosting many interacting modes[23, 24, 25, 26] offer promising neuromorphic platforms. Their lasing thresholds, multiple diverse modes, and competition for optical gain between modes[27, 26, 28, 29] provide intrinsic heterogenous nonlinearity, missing or weak in many photonic schemes[3, 30, 13, 2], while their sensitivity to structured optical input provides spatial processing[27, 29, 26, 31]. The mode competition inhibitively couples modes, as they antagonistically try to harvest a finite amount of pump energy. Despite these properties, such systems have remained unexplored for computation.
Here, we present a retinomorphic machine vision system implemented in a 150 m diameter on-chip semiconductor random network laser[32], fabricated through the mature process of wafer-bonding InP on oxide-coated Si, followed by electron beam lithography and reactive ion etching[33]. The large number of spatially-overlapping lasing modes provide effective heterogenous photonic neurons, performing parallel image feature detection through their spatially-sensitive response, with nonlinearity arising from competition for optical gain between neighbouring coupled modes - a photonic analogue of retinal inhibition. Beyond feature extraction, we demonstrate higher-level vision tasks of classification and segmentation, with strong performance in few-shot/low-data regimes. Crucially, the system outperforms a range of software neural networks including EfficientNet, a large-scale modern CNN, on hard biomedical tasks with scarce and class-imbalanced training data.
These results position the system not just as a novel photonic device addressing the size and nonlinearity issues facing optical computing, but as an exemplar of a new class of compact, few-shot capable neuromorphic hardware. By embedding bio-inspired behaviour combining excitation, inhibition, and nonlinear neuronal heterogeneity directly into the physical dynamics, we enable neuromorphic processing that meets the demands of edge AI.
@@
Our physical system is a 150 m diameter InP network laser comprising interconnected lateral nanoscale waveguides, lithographically patterned and etched from a Si-bonded InP layer[32, 33]. The network has a random Voronoi topology with three waveguides meeting at each vertex (average waveguide length = 5 m). Fabrication and optical setup details are provided in the Methods and supplementary information. A pump beam, spatially-structured via digital micromirror device (DMD) (figure 1a), induces optical gain in the InP network (fig. 1b). Photons are generated by spontaneous emission and scatter through the waveguides, amplified by stimulated emission from the illuminated InP. Constructive interference along specific scattering paths defines optical modes, which undergo net gain or loss depending on the extent of overlap between the illumination pattern and the mode profiles. The large number of possible scattering paths results in many spatially-distributed overlapping modes which nonlinearly compete for gain. Pump illumination is supplied by a 633 nm, 200 fs pulse laser, with pulse energy measured at 43 nJ at the sample (under full illumination). The spatial distribution of light emitted from the network by three different modes is illustrated by the coloured network nodes in figure 1b), experimentally imaged by a frequency-resolved slit-scan camera. The network lasing response is measured using a two-dimensional CCD detector with diffraction grating. This allows collection of two dimensional hyperspectra (fig. 1c) with one spectral dimension, and one spatial dimension at 1m spatial resolution. Coloured boxes on the hyperspectrum correspond to the spectral location of six feature-detecting modes examined in fig. 1 f-i) (additional details in supplementary information).
@@
Figure 1d) shows a schematic of feature detection in conventional software CNNs. A kernel filter is raster-scanned across small image ‘windows’ of an input image. The kernel values are matrix multiplied against image pixel values, and the results summed to a single number which becomes the pixel value of an output feature map, centered at the image window location. This process is repeated across the input image to build feature maps, with different kernel filters sensitive to different image features e.g. left or right edges. To detect multiple image features, the process is repeated sequentially using different kernels - consuming time and energy. This process is entirely linear, with subsequent nonlinear activations performed at time and energy cost in order to use the feature maps for higher-level AI tasks such as image classification or segmentation. This linear approach is what has so-far been employed in physical convolutional feature detection schemes[34, 35, 36, 37].
@@
Figure 1e) shows a schematic of the network laser feature detection. Image windows are illuminated on the network in a raster-scan fashion similar to CNNs, but here the network simultaneously detects many features in parallel via multiple lasing modes, with integrated physical nonlinearity. One scan produces many feature maps in parallel, ten experimentally (forty in simulation). Our network occupies a compact 150 m footprint, relative to the mm/cm-scale of alternative physical feature-detecting schemes.
@@
Figure 1f-i) show experimental parallel feature detection results for a ‘Cameraman’ test image (f). Example image windows (4 pixels, inset of f) are shown in (g) alongside their respective lasing hyperspectra (2nd col.); LL plots (3rd col., network lasing mode intensity vs pump intensity) which show strongly heterogenous nonlinear responses for each mode, and a variation in these nonlinearities depending on the input image; and a histogram of mode intensities (4th col.) for six feature-detecting modes, taken at the pump power indicated by dashed vertical lines in the LL plots. Mode lasing intensity is increased or suppressed by the presence of different image features, due to a combination of correlation between the input image shape and spatial mode profile, and nonlinear mode competition between spatially-overlapping modes. This enables feature maps to be obtained by plotting the lasing amplitude of different modes above the uniform illumination level at different positions in the raster-scan. The mode intensities plotted in the histogram in (g) are used to generate the pixel amplitudes of our feature maps. Figure 1h) shows a composite feature map produced by six lasing modes, with different colours corresponding to the lasing amplitude of six modes indicated in fig. 1c,g-i). Figure 1j) shows separate feature maps for the modes comprising the composite map. Feature maps for all ten feature-detecting modes are shown in supplementary information.
@@
a) 50 m network topology used for netSALT simulations a-f). b) Three pump illumination patterns: uniform, horizontal edge (‘right’) and vertical edge (‘top’). c) Simulated network lasing spectra for the three illumination patterns. Three modes A-C are indicated, corresponding to d). d) Simulated lasing intensities of three modes A-C, for the three pump illumination patterns. Mode competition nonlinearly suppresses specific modes below lasing threshold, and amplifies other modes dependent on input pump image. Spatial mode profiles are shown inset in left panel. e) Input image, ‘CL’ for ‘Complex Lasing’. The input image is fractured into 88 pixel windows and illuminated onto the network sequentially in a raster-scan fashion, similar to kernel scanning in a convolutional filter process. f) Simulated feature maps for five modes. Pixel brightness in the feature maps corresponds to mode lasing intensity for that position in the window raster-scan. Different modes lase in response to different image features, spatial mode profiles shown inset. A larger set of 28 simulated feature maps and corresponding mode profiles is shown in supplementary information. g) Experimentally measured ‘CL’ feature maps produced from six modes, with different lasing modes responding to specific image features. h) Experimental composite feature map combining modes from g). i) Equivalent 44 kernels for the six lasing modes in (g,h), extracted by linear fits to experimental data. Coloured frames correspond to mode feature map colour scheme. j,k) Colour photograph of a mandrill. Image is separated into RGB channels and feature detection is performed at multiple kernel sizes demonstrating the resolution flexibility of our scheme. Composite RGB feature maps are shown in (k) for 44, 77 and 1111 kernels illustrating the range of high-to-low spatial frequency features detected as kernel size is increased.
@@
We can use simulations of the photonic physics to examine the origin of the feature-detection. Here, we employ netSALT[26], a numerical model which solves the nonlinear interaction of optical waves on graph networks within the steady-state ab-initio laser theory (SALT)[38] approximation. The netSALT model includes amplification and loss on graph edges and nonlinear competition between spatially-overlapping lasing modes. Figure 2a) shows a random Voronoi graph network used for the netSALT simulations. Network topology and average edge length (5 m) match the experimentally measured network shown in fig. 1, with a smaller 50 m diameter used in simulation to reduce computational load.
Three illumination patterns are explored (fig. 2b): uniform illumination (‘full’), horizontal edge (‘right’) and vertical edge (‘top’). The resultant simulated lasing spectra are shown in figure 2c), with strong variation in lasing spectra under the different illumination patterns.
NetSALT predicts 239 possible lasing modes for the network considered here, with 172 reaching lasing threshold for at least one of the illumination patterns. The three spectra in fig. 2c) show the sum of all active modes for the different illumination patterns. We focus on three specific modes, labelled A,B,C, identified in figure 2c) by dashed vertical lines. Figure 2d) shows the lasing amplitude of the three modes under ‘full’, ‘right’ and ‘top’ illumination patterns with the spatial mode profiles inset in the left-hand panel. The modes are spatially-distributed across the network and overlap considerably. As multiple modes are active in the same network waveguides, nonlinear mode-competition occurs between the modes as they antagonistically compete for finite pump gain. Under full illumination, mode B shows the highest amplitude, while mode C does not lase as it falls below threshold. For a right-edge illumination, mode B now falls below the lasing threshold and is deactivated, allowing mode C to lase with the highest amplitude as it no longer faces strong mode-competition from the mode B. For a top-edge illumination, mode A exhibits far higher amplitude and mode C now falls below the lasing threshold.
@@
We now consider the netSALT-simulated feature-detection performance of the system. An input image of the letters CL (‘Complex Lasing’) is shown in fig. 2e). The image was fractured into 8x8 pixel windows and illuminated on the network in a raster-scan fashion. The lasing response was simulated for each window illumination, and the mode lasing amplitude plotted at each position in the raster scan to generate separate feature-maps for each mode. Figure 2f) shows feature maps provided by five different modes, with the corresponding spatial mode profiles inset. These plots demonstrate the ability of the network to perform spectrally-multiplexed feature-detection of a range of image features, with the broad set of lasing modes effectively providing parallel convolutional filters. The spatial location of the mode profiles within the network are seen to show correspondence with the detected image features, e.g. modes detecting lower image edges are located towards the bottom of the network (2nd panel), modes detecting left edges are located towards the left of the network (4th panel), and so on. Of the 172 active modes in the simulated network, 40 were identified as exhibiting good feature-detection. A set of 28 selected feature maps and corresponding spatial mode profiles is shown in the supplementary information, alongside the full set of 172 feature maps from all modes in supplementary information. A regular hexagonal network was simulated to compare against the random topology (supplementary information), with far fewer detected features in the hexagonal network. This demonstrates the ability of network topology to control the device functionality, with further explorations into the correspondence between topology and performance providing fruitful ground for future work.
@@
g,h,i) Classification accuracy vs. training set length, comparing single- and multi-layer photonic network accuracy against software benchmarks: logistic regression, multilayer perceptron (MLP), LeNet-5 convolutional neural network (CNN), a larger CNN (96 filters), and the far larger EfficientNetV2-B0 CNN (3040 filters, in a transfer-learning scheme with pre-trained ImageNet weights). Maximum accuracies for the multi-layer photonic network of Digits: 98.05%, Fashion: 87.85%, and BreaKHis: 90.12% are observed. Zoomed-in regions of the few-shot regime are shown in (g,h), where the photonic network exceeds software benchmark performance.
@@
We now examine experimental feature detection. Figure 2g) shows six experimentally measured feature maps of the ‘CL’ image with colours corresponding to the modes shown in figure 1 and a composite feature map shown in figure 2h). Parallel feature detection of multiple image features is observed experimentally, with close correspondence to the netSALT simulated behaviour. Supplementary figures examine the feature detection performed by ten lasing modes across a range of test images. To compare our physically-detected feature maps against linear matrix kernel filters as used in software CNNs, we extract effective kernel filters for our modes via least-squares fitting between input images and the experimentally measured mode intensity (details in supplementary information). The extracted effective filters are shown in fig. 2i) with further details in supplementary information. Notably, both positive and negative kernel weights are observed - generating bipolar weights is highly challenging in neuromorphic hardware. This is an attractive benefit of the photonic lateral inhibition processes in our system, which enable mode amplitudes to both increase and decrease in response to specific image features. The values below each effective kernel indicate how well the linear fitting process captures our experimental results. When performing the same kernel fitting process on arbitrary linear software kernel filters, the process performs perfectly with values of 1.0. The kernels extracted from our experimental data have values from 0.628-0.946 (see SI), suggesting that aspects of the physical feature detection process cannot be fully captured by linear kernel filters.
@@
We extend experimental demonstration of feature detection to high-resolution colour images, showing a flexible range of kernel sizes using the popular ‘mandrill’ test image (512512 pixels, not a baboon as often thought!)[39]. Figure 2j) shows the input image, which is separated into RGB channels and illuminated onto the network at a range of kernel sizes to demonstrate the capability of our system to detect features with arbitrary kernel window size. This flexibility is typically not available in the more common matrix-multiplication style feature detection schemes where kernel size is hard-coded at fabrication stage[34]. Figure 2k) shows experimentally detected feature maps for 44, 77 and 1111 kernels, with the R/G/B coloured features detected from the corresponding colour channel (the response of all ten feature-detecting modes are summed to produce a single feature map for each colour channel here). The ability to tune feature detection sensitivity to higher or lower spatial frequency is shown in our feature maps produced by different kernel sizes, the 44 kernel picks up high-frequency details in the mandrill fur around image corners, whereas the 1111 kernel exhibits weaker response to these. Increasingly, powerful modern software CNNs such as EfficientNet[40, 41] employ a range of filter sizes, demonstrating the utility and attraction of our ability to freely select filter size. Separate lasing mode feature maps are shown for a wider range of test images in supplementary information.
@@
We perform classification on three tasks: MNIST digits, Fashion-MNIST, and the challenging low-data BreaKHis breast cancer diagnosis[42] (colour histopathology images, 400 magnification, 700460 resolution). Figures 3d,e,f) show example images for each task, with coloured lines illustrating feature maps detected by the network. For MNIST digits and Fashion, we experimentally process 40,000 training images with a test set of 4000 images and training sets of 125 to 36,000 images. For BreaKHis we process 1693 images, with a test set of 250 and training sets of 10 to 1,443 images. BreaKHis contains a roughly 2:1 ratio of malignant to benign images, with this class-imbalance increasing task difficulty.

### v5

# Few-Shot Neuromorphic Vision in a Nonlinear Photonic Network Laser

@@ Abstract
With the growing prevalence of AI, demand increases for hardware that mimics the brain’s ability to extract structure from limited data. In the retina, ganglion cells detect features from sparse inputs via lateral inhibition, where neurons antagonistically suppress activity of neighbouring cells. Biological neurons exhibit diverse heterogeneous nonlinear responses, linked to robust learning and strong performance in low-data regimes.
Here, we introduce a retinally-inspired photonic computing system where spatially-competing lasing modes in a random network laser act as heterogeneous, inhibitively-coupled neurons - enabling feature detection, few-shot classification, and segmentation.
This silicon-compatible scheme harnesses heterogeneous excitatory and inhibitory nonlinear physical dynamics which give rise to emergent photonic computing behaviour, including parallel feature detection and strong performance when training data is scarce. We report 98.05% and 87.85% accuracy on MNIST and Fashion-MNIST, and 90.12% on BreaKHis cancer diagnosis - outperforming software CNNs including EfficientNetV2 and the vision transformer ViT in few-shot and class-imbalanced regimes with training sets of up to several hundred images. We demonstrate combined segmentation and classification on the HAM10k skin lesion dataset, achieving DICE and Jaccard scores of 84.49% and 74.80%. These results demonstrate the potential of random lasing networks as nonlinear photonic learning systems, and highlight the ability of heterogeneous nonlinear dynamics to support strong learning in challenging low-data scenarios.
@@ Introduction
Physics-based neuromorphic computing systems[1, 12, 3, 13, 14, 15, 16, 17] have been explored as candidates for edge-AI hardware[14], but integrating the above qualities to achieve software-equivalent performance on harder problems with very limited training data, such as biomedical tasks, remains challenging.
@@
Biological systems offer a compelling blueprint. In the retina, ganglion cells perform spatially-sensitive nonlinear processing via lateral inhibition[18, 19, 20, 21], where neurons compete to fire while inhibitively suppressing the activity of neighbouring cells. This mechanism enables efficient encoding of spatial structure, even under sparse or noisy input, and the ability to detect many features in parallel from multiple ganglion cells[18, 22]. The combination of excitatory dynamics - firing in response to input, and inhibitory ones - suppressing activity of nearby coupled neurons, is crucial to the strong performance of biological learning systems, but to date physics-based neuromorphic schemes typically focus solely on engineering excitatory dynamics. Addressing this inhibition is an inviting route towards taking neuromorphic performance closer to the biological systems which inspire it. In addition, the diverse heterogeneous nonlinear responses across ensembles of biological neurons have been linked to robust learning[23], and creating effective high-dimensional latent spaces which have been shown to give rise to strong few-shot learning in software networks[11].
Photonic computing schemes have demonstrated promising routes to high-bandwidth, low-energy performance, with the realisation of efficient optical nonlinearities identified as a key area requiring further development[2, 24, 25]. Recent studies have shown that access to strong nonlinearities can substantially enhance computational capability, for example through opto-electronic nonlinear materials[26], nonlinear mode mixing in multi-mode fibres[17], or nonlinear dynamics such as lasing mode competition in multi-mode VCSELs[27, 28]. Exploration of combining heterogeneous nonlinearities is particularly promising here, integrating multiple distinct nonlinear dynamics - which has been shown to give rise to robust learning in software models of biologically-inspired neurons[23]. Developing physical learning systems which exhibit both strong heterogeneous nonlinearities and the capability for useful feature extraction is hence highly attractive, with experimental demonstration of schemes integrating these qualities so-far elusive.
Here, we present a retinally-inspired neuromorphic vision system exhibiting strong heterogeneous photonic nonlinearities, parallel feature detection, and strong learning performance in the few-shot/low-data regimes where training examples are scarce. Our system is implemented in a 150 m diameter on-chip semiconductor random network laser[29], fabricated through the mature process of wafer-bonding InP on oxide-coated Si, followed by electron beam lithography and reactive ion etching[30].
These complex random lasers host many strongly interacting lasing modes[31, 32, 33, 34], with nonlinear physics well-suited for use in a neuromorphic system. The large population of modes is spatially and spectrally varying with complex disordered distributions, and provides strong intrinsic heterogeneous photonic nonlinearity - including both excitatory dynamics, where modes reach their threshold and lasing activates, and inhibitory dynamics, where spatially-overlapping modes antagonistically compete for finite optical gain (termed ‘mode competition’)[35, 34, 36, 37]. This combination of competing processes mirrors the coupled excitatory and inhibitory dynamics of retinal ganglion cells, where cells attempting to activate in response to incoming light are laterally inhibited by neurotransmitters from neighbouring cells. In the retina, this directly gives rise to biological feature and edge detection[22, 21, 20, 19, 18].
The networks’ sensitivity to structured optical input[35, 37, 34, 38] makes them an intrinsically well-matched platform for spatial processing. We demonstrate that the competing excitatory and inhibitory nonlinear photonic dynamics give rise to parallel image feature detection, in addition to the higher-level downstream vision tasks of classification and segmentation – with strong performance in few-shot/low-data regimes. Crucially, the system outperforms a range of software neural networks when training data is scarce, including a pre-trained vision transformer[39] (ViT_b_16, 86 million parameters) and a pre-trained large-scale CNN[39] (EfficientNetV2-B0, 7.9 million parameters) on hard biomedical tasks with scarce and class-imbalanced training data.
These results motivate the integration of heterogeneous nonlinear physical dynamics and feature extraction capabilities in physics-based learning systems as a highly-promising direction for performing challenging few-shot/low-data learning tasks, crucial to deliver the growing demand for edge-computing systems capable of performing both inference and training[40, 41]. Additionally, we show that random network lasers can provide such integrated capability in on-chip photonic systems. By embedding bio-inspired behaviour combining excitation, inhibition, and nonlinear neuronal heterogeneity directly into the physical dynamics, we enable neuromorphic processing that meets the demands of edge AI.
@@
Our physical system is a 150 m diameter InP network laser comprising interconnected lateral nanoscale waveguides, lithographically patterned and etched from a Si-bonded InP layer[29, 30]. The network has a random Voronoi topology with three waveguides meeting at each vertex (average waveguide length = 5 m). Fabrication and optical setup details are provided in the Methods and supplementary figure S1. A pump beam, spatially-structured via digital micromirror device (DMD) (figure 1a), induces optical gain in the InP network (fig. 1b). Photons are generated by spontaneous emission and scatter through the waveguides, amplified by stimulated emission from the illuminated InP. Constructive interference along specific scattering paths defines optical modes, with the vast number of potential scattering paths giving rise to a large number of spatially-distributed distinct overlapping lasing modes which nonlinearly compete for gain. The modes undergo net gain or loss depending on the extent of overlap between the illumination pattern and the mode profiles. The 100 nm bandwidth of the lasing spectra is given by the gain spectrum of the network material (InP), roughly centred around the bandgap.
Pump illumination is supplied by a 633 nm, 200 fs pulse laser, with pulse energy measured at 43 nJ at the sample (under full illumination). The spatial distribution of light emitted from the network by three different modes is illustrated by the coloured network nodes in figure 1b), experimentally imaged by a frequency-resolved slit-scan camera. The network lasing response is measured using a two-dimensional CCD detector with diffraction grating. This allows collection of two dimensional hyperspectra (fig. 1c) with one spectral dimension, and one spatial dimension at 1m spatial resolution, and each hyperspectrum is acquired by a single camera image acquisition. The broad lasing modes in the spectral dimension originate from band-filling and rapid carrier-induced index changes caused by ultrashort pulse excitation. Coloured boxes on the hyperspectrum correspond to the location of six feature-detecting modes examined in figure 1 f-i) (additional details in supplementary figure S2).
@@
Figure 1d) shows a schematic of feature detection in conventional software CNNs. A kernel filter is raster-scanned across small image ‘windows’ of an input image. The kernel values are matrix multiplied against image pixel values, and the results summed to a single number which becomes the pixel value of an output feature map, centered at the image window location. This process is repeated across the input image to build feature maps, with different kernel filters sensitive to different image features e.g. left or right edges. To detect multiple image features, the process is repeated sequentially using different kernels - consuming time and energy. This process is entirely linear, with subsequent nonlinear activations performed at time and energy cost in order to use the feature maps for higher-level AI tasks such as image classification or segmentation. This linear matrix multiplication approach is what has so-far been employed in physical convolutional feature detection schemes[42, 43, 44, 45].
@@
Figure 1e) shows a schematic of the network laser feature detection. Image windows are illuminated on the network in a raster-scan fashion similar to CNNs, with the hyperspectral channels of the network’s lasing emission acting as the output. While input parallelisation through input wavelength multiplexing has been achieved in photonic schemes[42], we achieve feature detection parallelisation via the output spectral channels, where different modes at different wavelengths and spatial positions provide sensitivity to different image features. One raster-scan produces multiple feature maps, with integrated physical nonlinearity, ten experimentally and forty in simulation. Our network occupies a compact 150 m footprint on chip, alongside other off-chip optical components (e.g., laser source, DMD, and camera, see supplementary information).
Figure 1f-i) show experimental parallel feature detection results for a ‘Cameraman’ test image (f). Example image windows (4 pixels, inset of f) are shown in (g) alongside their respective lasing hyperspectra (2nd col.), light-in versus light-out (LL) plots (3rd col., network lasing mode intensity vs pump intensity), and a histogram of mode intensities (4th col.) for six feature-detecting modes, taken at the pump power indicated by dashed vertical lines in the LL plots.
Mode lasing intensity is increased or inhibited by the presence of different image features, due to a combination of correlation between the input image shape and spatial mode profile, and nonlinear mode competition between spatially overlapping modes. This above-threshold mode competition is intrinsic to multimode lasers. As overlapping modes compete for gain, they inhibit each other by spatial and spectral ‘hole burning’ effects[46, 47]. These effects can be modelled using steady-state ab-initio laser theory (SALT)[48], which we have previously used to build a graph-based theoretical framework of network lasers - termed ‘netSALT’[34]. The large variety of spatially and spectrally varying modes in our random network laser gives rise to strong mode competition, and high sensitivity to pump patterns (e.g. input images)[34, 29]. The resulting increase or inhibition of lasing modes with increasing pump power is clearly seen in the LL-plots in figure1g), which show strongly heterogeneous nonlinear responses for each mode, which vary substantially depending on the input image.
@@
This combination of excitatory and inhibitory nonlinear responses to input images enables feature maps to be obtained by plotting the lasing amplitude of different modes above the uniform illumination level at different positions in the raster-scan. The mode intensities plotted in the histogram in (g) are used to generate the pixel amplitudes of our feature maps. Figure 1h) shows a composite feature map produced by six lasing modes, with different colours corresponding to the lasing amplitude of six modes indicated in fig. 1c,g-i). Figure 1i) shows separate feature maps for the modes comprising the composite map. See Methods for raster-scan details. Feature maps for all ten feature-detecting modes are shown in supplementary figure S11.
@@
a) 50 m network topology used for netSALT simulations a-f). b) Three pump illumination patterns: uniform, horizontal edge (‘right’) and vertical edge (‘top’). c) Simulated network lasing spectra for the three illumination patterns. Three modes A-C are indicated, corresponding to d). d) Simulated lasing intensities of three modes A-C, for the three pump illumination patterns. Mode competition nonlinearly suppresses specific modes below lasing threshold, and amplifies other modes dependent on input pump image. Spatial mode profiles are shown inset in left panel. e) Input image, ‘CL’ for ‘Complex Lasing’. The input image is fractured into 88 pixel windows and illuminated onto the network sequentially in a raster-scan fashion, similar to kernel scanning in a convolutional filter process. f) Simulated feature maps for five modes. Pixel brightness in the feature maps corresponds to mode lasing intensity for that position in the window raster-scan. Different modes lase in response to different image features, spatial mode profiles shown inset. A larger set of 28 simulated feature maps and corresponding mode profiles is shown in supplementary information. g) Experimentally measured ‘CL’ feature maps produced from six modes, with different lasing modes responding to specific image features. h) Experimental composite feature map combining modes from g). i) Equivalent 44 kernels for the six lasing modes in (g,h), extracted by linear fits to experimental data. Coloured frames correspond to mode feature map colour scheme. j,k) Colour photograph of a mandrill. Image is separated into RGB channels and feature detection is performed at multiple kernel sizes demonstrating the resolution flexibility of our scheme. Composite RGB feature maps are shown in (k) for 44, 77 and 1111 kernels illustrating the range of high-to-low spatial frequency features detected as kernel size is increased. The sizes of each kernel are illustrated below the respective maps.
@@
We can use simulations of the photonic physics to examine the origin of the feature-detection. Here, we employ netSALT[34], a numerical model which solves the nonlinear interaction of optical waves on graph networks within the steady-state ab-initio laser theory (SALT)[48] approximation. The netSALT model includes amplification and loss on graph edges and nonlinear competition between spatially-overlapping lasing modes. Figure 2a) shows a random Voronoi graph network used for the netSALT simulations. Network topology and average edge length (5 m) match the experimentally measured network shown in fig. 1, with a smaller 50 m diameter used in simulation to reduce computational load. Due to the smaller network size (total edge length), the simulated network will have fewer modes than the experimental one.
Three illumination patterns are explored (fig. 2b): uniform illumination (‘full’), horizontal edge (‘right’) and vertical edge (‘top’). The resultant simulated lasing spectra are shown in figure 2c), with strong variation in lasing spectra under the different illumination patterns. NetSALT predicts 239 possible lasing modes for the network considered here, with 172 reaching lasing threshold for at least one of the illumination patterns. The three spectra in fig. 2c) show the sum of all active modes for the different illumination patterns. We focus on three specific modes, labelled A,B,C, identified in figure 2c) by dashed vertical lines. Figure 2d) shows the lasing amplitude of the three modes under ‘full’, ‘right’ and ‘top’ illumination patterns, with the spatial mode profiles inset in the left-hand panel. The modes are spatially-distributed across the network and overlap considerably. As multiple modes are active in the same network waveguides, nonlinear mode-competition occurs between the modes as they antagonistically compete for finite pump gain. Under full illumination, mode B shows the highest amplitude, while mode C does not lase as it falls below threshold. For a right-edge illumination, mode B now falls below the lasing threshold and is deactivated, allowing mode C to lase with the highest amplitude as it no longer faces strong mode-competition from mode B. For a top-edge illumination, mode A exhibits far higher amplitude and mode C now falls below the lasing threshold.
@@
We now consider the netSALT-simulated feature-detection performance of the system. An input image of the letters CL (‘Complex Lasing’) is shown in fig. 2e). The image was fractured into 8x8 pixel windows and illuminated on the network in a raster-scan fashion. The lasing response was simulated for each window illumination, and the simulated mode lasing amplitude plotted at each position in the raster scan to generate separate feature-maps for each mode. Figure 2f) shows feature maps provided by five different modes, with the corresponding spatial mode profiles inset. These plots demonstrate the ability of the network to perform spectrally-multiplexed feature-detection of a range of image features, with the broad set of lasing modes effectively providing parallel convolutional filters. The spatial location of the mode profiles within the network are seen to show correspondence with the detected image features, e.g. modes detecting lower image edges are located towards the bottom of the network (2nd panel), modes detecting left edges are located towards the left of the network (4th panel), and so on. Of the 172 active modes in the simulated network, 40 were identified by inspection as exhibiting good feature-detection. A set of 28 selected feature maps and corresponding spatial mode profiles is shown in the supplementary figure S4, alongside the full set of 172 feature maps from all modes in supplementary figure S5. A regular hexagonal network was simulated to compare against the random topology (supplementary figure S7), with far fewer detected features in the hexagonal network. This demonstrates the ability of network topology to control the device functionality, with further explorations into the correspondence between topology and performance providing fruitful ground for future work.
@@
g,h,i) Classification accuracy vs. training set length, comparing single- and multi-layer photonic network accuracy against software benchmarks: logistic regression, multilayer perceptron (MLP), LeNet-5 convolutional neural network (CNN), a 96-filter CNN, EfficientNetV2-B0 CNN (3040 filters, in a transfer-learning scheme with pre-trained ImageNet weights) , and a vision transformer (ViT) with 86 million parameters. Maximum accuracies for the multi-layer photonic network of Digits: 98.05%, Fashion: 87.85%, and BreaKHis: 90.12% are observed. Zoomed-in regions of the few-shot regime are shown in (g,h), where the photonic network exceeds software benchmark performance.
@@
We now examine experimental feature detection. Figure 2g) shows six experimentally measured feature maps of the ‘CL’ image with colours corresponding to the modes shown in figure 1 and a composite feature map shown in figure 2h). Parallel feature detection of multiple image features is observed experimentally, with close correspondence to the netSALT simulated behaviour. Supplementary figures S8, S9 and S10 examine the feature detection performed by ten lasing modes across a range of test images. To compare our physically-detected feature maps against linear matrix kernel filters as used in software CNNs, we extract effective kernel filters for our modes via least-squares fitting between input images and the experimentally measured mode intensity (details in supplementary information). The extracted effective filters are shown in fig. 2i) with further details in supplementary fig. S20. Notably, both positive and negative kernel weights are observed - generating bipolar weights is highly challenging in neuromorphic hardware. This is an attractive benefit of the photonic lateral inhibition processes in our system, which enable mode amplitudes to both increase and decrease in response to specific image features. The values below each effective kernel indicate how well the linear fitting process captures our experimental results. When performing the same kernel fitting process on arbitrary linear software kernel filters, the process performs perfectly with values of 1.0. The kernels extracted from our experimental data have values from 0.628-0.946 (see SI fig. S20 and S21), suggesting that aspects of the physical feature detection process cannot be fully captured by linear kernel filters. This is further supported by a similar linear-fit study on netSALT-simulated data shown in SI fig. S22. The complexity of the fitted kernels also indicates that while these modes were chosen by inspection to produce edge-detecting feature maps, the underlying dynamics are likely to support other convolutional filters.
We extend experimental demonstration of feature detection to high-resolution colour images, showing a flexible range of kernel sizes using the popular ‘mandrill’ test image (256 246 pixels, not a baboon as often thought!)[49, 50]. Figure 2j) shows the input image, which is separated into RGB channels and illuminated onto the network at a range of kernel sizes to demonstrate the capability of our system to detect features with arbitrary kernel window size. This shows a flexibility of our system in handling input images and kernel windows of varying resolution.
@@
Figure 2k) shows experimentally detected feature maps for 44, 77 and 1111 kernels, with the R/G/B coloured features detected from the corresponding colour channel (the response of all ten feature-detecting modes are summed to produce a single feature map for each colour channel here; the separate mode maps can be seen in Supplementary Figures S17, S18, and S19). The ability to tune feature detection sensitivity to higher or lower spatial frequency is shown in our feature maps produced by different kernel sizes, the 44 kernel picks up high-frequency details in the mandrill fur around image corners, whereas the 1111 kernel exhibits weaker response to these. Increasingly, powerful modern software CNNs such as EfficientNet[51, 52] employ a range of filter sizes, demonstrating the utility and attraction of our ability to freely select filter size. Separate lasing mode feature maps are shown for a wider range of test images in supplementary figures S11, S12 and S14.
@@
We perform classification on three tasks: MNIST digits, Fashion-MNIST, and the challenging low-data BreaKHis breast cancer diagnosis[53] (colour histopathology images, 400 magnification, 700460 resolution). Figures 3d,e,f) show example images for each task, with coloured lines illustrating feature maps detected by the network. For MNIST digits and Fashion, we experimentally process 40,000 training images with a test set of 4000 images and training sets of 125 to 36,000 images. For BreaKHis we process 1693 images, with a test set of 250 and training sets of 10 to 1,443 images. BreaKHis contains a roughly 2:1 ratio of malignant to benign images, with this class-imbalance increasing task difficulty. No augmentation is performed on the original datasets to synthetically increase their length.